# Chaos propagation in genetic algorithms: An optimal transport approach

### v1

@@ 3.4 Limit as $\tau\to 0$ and final estimate
Now, we show that the explicit time discretization $f_{n}$ (3.1) converges to the solution of the PDE $f(t_{n})$ (1.2) as $\tau\to 0$ in the norm $\|f_{n}-f(t_{n})\|_{\mathtt{KR}}$. We expect an explicit Euler discretization to return an error of order $\mathcal{O}(\tau)$. The proof is standard, but we give a brief sketch for completeness.
We note that the RHS of the Boltzmann-like equation is Lipschitz with respect to $\|\cdot\|_{\mathtt{KR}}$ since, thanks to Lemma 2.3, for $f,g\in\mathcal{P}_{q}(\mathbb{R}^{d})$, $q\geq 1$
$\langle\phi,Q^{+}(f,f)-Q^{+}(g,g)\rangle\leq\kappa_{\mathbf{F}}\left|\langle|\phi|,f-g\rangle\right|\leq\kappa_{\mathbf{F}}\sup_{\textup{Lip}(\psi)\leq 1}\langle\psi,f-g\rangle\,.$
for any test function $\phi,\textup{Lip}(\phi)\leq 1$, and in particular the one realizing $\|Q^{+}(f,f)-Q^{+}(g,g)\|_{\mathtt{KR}}$.
Moreover, using the moment bounds on $f$ (Lemma 3.2), one obtains the uniform estimate
$\|f(t)-f(s)\|_{\mathtt{KR}}\leq C\,(1+\kappa_{\mathbf{F}})\,(M_{q}(f_{0})^{1/q}+\sigma)\,e^{\tilde{C}\,\kappa_{\mathbf{F}}T}\,|t-s|.$
Let $e_{n}:=\|f_{n}-f(t_{n})\|_{\mathtt{KR}}$. First, we remark that the weak formulation in Definition 2.1 also holds for Lipschitz functions, since $f(t)\in\mathcal{P}_{q}(\mathbb{R}^{d})$ for $t\in[0,T]$. This can be shown by approximating $\phi$ with $\phi_{R}=\max(-R,\min(\phi,R))\in C_{b}(\mathbb{R}^{d})$ and eventually taking the limit $R\to\infty$. By the weak formulation of (1.2), we have
@@
Taking the supremum over $\phi$ with $\|\phi\|_{\textup{Lip}}\leq 1$ and $\phi\geq 0$, and using the Lipschitz stability of $Q^{+}$ in $\|\cdot\|_{\mathtt{KR}}$, we deduce
$e_{n+1}$ $\leq(1-\tau)e_{n}+\tau\kappa_{\mathbf{F}}e_{n}+\int_{t_{n}}^{t_{n+1}}\Big(\kappa_{\mathbf{F}}\|f(t_{n})-f(s)\|_{\mathtt{KR}}+\|f(s)-f(t_{n})\|_{\mathtt{KR}}\Big)\,\mathrm{d}s$
$\leq\big(1+\tau(\kappa_{\mathbf{F}}-1)\big)e_{n}+(1+\kappa_{\mathbf{F}})\int_{t_{n}}^{t_{n+1}}\|f(s)-f(t_{n})\|_{\mathtt{KR}}\,\mathrm{d}s$
$\leq\big(1+\tau(\kappa_{\mathbf{F}}-1)\big)e_{n}+(1+\kappa_{\mathbf{F}})\int_{t_{n}}^{t_{n+1}}\tilde{C}\,(1+\kappa_{\mathbf{F}})\,e^{\tilde{C}\,\kappa_{\mathbf{F}}T}(s-t_{n})\,\mathrm{d}s$
$\leq\big(1+\tau(\kappa_{\mathbf{F}}-1)\big)e_{n}+\tilde{C}\,(1+\kappa_{\mathbf{F}})^{2}\,e^{\tilde{C}^{\prime}\,\kappa_{\mathbf{F}}T}\tau^{2}\,$,
@@
Iterating the above inequality for $t_{n}=n\tau\leq T$ and using $e_{0}=0$ gives for some other constants $C,C^{\prime}>0$ (depending on the same arguments)
$\sup_{t_{n}\leq T}\|f_{n}-f(t_{n})\|_{\mathtt{KR}}\leq C(1+\kappa_{\mathbf{F}}^{2})\tau e^{C^{\prime}\kappa_{\mathbf{F}}T},$ (3.7)
which shows that $\|f_{n}-f(t_{n})\|_{\mathtt{KR}}=\mathcal{O}(\tau)$ uniformly on $[0,T]$.
@@
$\sup_{t_{n}=n\tau\leq T}\|f^{N}_{n}-f(t_{n})\|_{\mathtt{KR}}$ $\leq\sup_{t_{n}=n\tau\leq T}\|f^{N}_{n}-f_{n}\|_{\mathtt{KR}}+\sup_{t_{n}=n\tau\leq T}\|f_{n}-f(t_{n})\|_{\mathtt{KR}}$
$\leq C(1+\kappa^{2}_{\mathbf{F}})Te^{C^{\prime}\kappa_{\mathbf{F}}T}\left(\,\varepsilon_{1}(N)\,+\,\tau\,\right)$
@@ Acknowledgment
GB was supported by the Wolfson Fellowship of the Royal Society “Uncertainty quantification, data-driven simulations and learning of multiscale complex systems governed by PDEs” of Prof. L. Pareschi at Heriot-Watt University.

### v2

@@ 3.4 Limit as $\tau\to 0$ and final estimate
Now, we show that the explicit time discretization $f_{n}$ (3.1) converges to the solution of the PDE $f(t_{n})$ (1.2) as $\tau\to 0$ in the BL norm $\|f_{n}-f(t_{n})\|_{\mathtt{BL}}$. We expect an explicit Euler discretization to return an error of order $\mathcal{O}(\tau)$. The proof is standard, see for instance [35], but we give a brief sketch for completeness.
We note that the RHS of the Boltzmann-like equation is Lipschitz with respect to $\|\cdot\|_{\mathtt{BL}}$ since, thanks to crossover-mutation stability (Lemma 2.3) and selection stability (Lemma 2.4), we have for any $f,g\in\mathcal{P}(\mathbb{R}^{d})$
$\|Q^{+}(f,f)-Q^{+}(g,g)\|_{\mathtt{BL}}\leq 2\left\|\frac{\mathbf{F}f}{\langle\mathbf{F},f\rangle}-\frac{\mathbf{F}g}{\langle\mathbf{F},g\rangle}\right\|_{\mathtt{BL}}\leq 2C_{\mathbf{F}}\|f-g\|_{\mathtt{BL}}\,.$
Using moment bounds for $f$ (similarly to the time-discrete result Lemma 3.2), it is possible to obtain a uniform estimate for the solution $f$ of the type
$\|f(t)-f(s)\|_{\mathtt{BL}}\leq C\,(M_{1}(f_{0})+\sigma)\,e^{\tilde{C}\,\kappa_{\mathbf{F}}T}\,|t-s|.$
Let $e_{n}:=\|f_{n}-f(t_{n})\|_{\mathtt{BL}}$. By the weak formulation of (1.2), we have for $\phi$, $\textup{Lip}(\phi)\leq 1$, $\|\phi\|_{\infty}\leq 1/2$
@@
Taking the supremum over $\phi$, and using the Lipschitz stability of $Q^{+}$ in $\|\cdot\|_{\mathtt{BL}}$, we deduce
$e_{n+1}$ $\leq(1-\tau)e_{n}+\tau 2C_{\mathbf{F}}e_{n}+\int_{t_{n}}^{t_{n+1}}\Big(2C_{\mathbf{F}}\|f(t_{n})-f(s)\|_{\mathtt{BL}}+\|f(s)-f(t_{n})\|_{\mathtt{BL}}\Big)\,\mathrm{d}s$
$\leq\big(1+\tau(2C_{\mathbf{F}}-1)\big)e_{n}+(1+2C_{\mathbf{F}})\int_{t_{n}}^{t_{n+1}}\|f(s)-f(t_{n})\|_{\mathtt{BL}}\,\mathrm{d}s$
$\leq\big(1+\tau(2C_{\mathbf{F}}-1)\big)e_{n}+(1+2C_{\mathbf{F}})\int_{t_{n}}^{t_{n+1}}\tilde{C}\,(1+2C_{\mathbf{F}})\,e^{\tilde{C}\,2C_{\mathbf{F}}T}(s-t_{n})\,\mathrm{d}s$
$\leq\big(1+\tau(2C_{\mathbf{F}}-1)\big)e_{n}+\tilde{C}\,(1+2C_{\mathbf{F}})^{2}\,e^{\tilde{C}^{\prime}\,2C_{\mathbf{F}}T}\tau^{2}\,$,
@@
Iterating the above inequality for $t_{n}=n\tau\leq T$ and using $e_{0}=0$ gives for some other constants $(\mathbf{F},q,M_{q}(f_{0})^{1/q},\sigma), C=C$$^{\prime}=C^{\prime}(\mathbf{F},q)>0$
$\sup_{t_{n}\leq T}\|f_{n}-f(t_{n})\|_{\mathtt{BL}}\leq C\tau e^{C^{\prime}\kappa_{\mathbf{F}}T},$ (3.7)
which shows that $\|f_{n}-f(t_{n})\|_{\mathtt{BL}}=\mathcal{O}(\tau)$ uniformly on $[0,T]$.
@@
$\sup_{t_{n}=n\tau\leq T}\|f^{N}_{n}-f(t_{n})\|_{\mathtt{BL}}$ $\leq\sup_{t_{n}=n\tau\leq T}\|f^{N}_{n}-f_{n}\|_{\mathtt{BL}}+\sup_{t_{n}=n\tau\leq T}\|f_{n}-f(t_{n})\|_{\mathtt{BL}}$
$\leq CTe^{C^{\prime}T}\left(\,\varepsilon_{1}(N)\,+\,\tau\,\right)$
@@ Acknowledgment
The author was supported by the Wolfson Fellowship of the Royal Society “Uncertainty quantification, data-driven simulations and learning of multiscale complex systems governed by PDEs” of Prof. L. Pareschi at Heriot-Watt University.